# On Central Primitives for Quantum Cryptography with Classical Communication

### v1

@@ Abstract
Recent work has introduced the “Quantum-Computation Classical-Communication” (QCCC) (Chung et. al.) setting for cryptography. There has been some evidence that One Way Puzzles ($\mathsf{OWPuzz}$) are the natural central cryptographic primitive for this setting (Khurana and Tomer). For a primitive to be considered central it should have several characteristics. It should be well behaved (which for this paper we will think of as having amplification, combiners, and universal constructions); it should be implied by a wide variety of other primitives; and it should be equivalent to some class of useful primitives. We present combiners, correctness and security amplification, and a universal construction for $\mathsf{OWPuzz}$. Our proof of security amplification uses a new and cleaner version construction of EFI from $\mathsf{OWPuzz}$ (in comparison to the result of Khurana and Tomer) that generalizes to weak $\mathsf{OWPuzz}$ and is the most technically involved section of the paper. It was previously known that $\mathsf{OWPuzz}$ are implied by other primitives of interest including commitments, symmetric key encryption, one way state generators ($\mathsf{OWSG}$), and therefore pseudorandom states ($\mathsf{PRS}$). However we are able to rule out $\mathsf{OWPuzz}$’s equivalence to many of these primitives by showing a black box separation between general $\mathsf{OWPuzz}$ and a restricted class of $\mathsf{OWPuzz}$ (those with efficient verification, which we call $\mathsf{EV-OWPuzz}$). We then show that $\mathsf{EV-OWPuzz}$ are also implied by most of these primitives, which separates them from $\mathsf{OWPuzz}$ as well. This separation also separates extending $\mathsf{PRS}$ from highly compressing $\mathsf{PRS}$ answering an open question of Ananth et. al.
@@ 1 Introduction
In the realm of cryptography, there is perhaps no primitive more important than the one-way function. A one-way function is an efficiently computable deterministic function which is easy to compute, but hard to invert. Although at first glance the definition seems simple, one-way functions are special for several reasons. First and foremost, one-way functions are “minimal.” If modern cryptography exists in any form, then one-way functions must also exist [HILL99, IL89, Imp95]. Furthermore, pretty much all of these constructions are obvious. Second, one-way functions are “useful.” There is a large class of cryptographic primitives (known as Minicrypt) which can all be built from and are equivalent to one-way functions [Imp95]. Included in Minicrypt are symmetric key encryption, pseudorandom generators, and commitment schemes [HILL99, GGM86, Nao91]. Finally, one-way functions are “well-behaved.” There satisfy several natural properties [Lev87], and are equivalent to most of their variants [Yao82, IL89]. Due to these three characteristics of one-way functions, one of the most useful things to do when trying to understand a new classical cryptographic primitive is to compare it to a one-way function.
@@
In the classical setting, it appears that one-way functions serve as an effective minimal primitive. In the quantum output setting, EFI pairs are a promising candidate for our minimal primitive. A number of recent works have also considered a hybrid setting, primitives where the cryptographic algorithms are quantum, but all communication and outputs are classical [${}^{+}$22, ALY23a]  [CLM23, KT24]. In the style of [${}^{+}$22], we will refer to this as the quantum computation classical communication (QCCC) setting. An immediate and natural question about this setting is “what is a good central primitive?”
@@
We also consider in detail an important variant of one-way puzzles, a restricted version of which was first introduced under the name “hard quantum planted problem for QCMA,” [KNY23], but which we will refer to as efficiently verifiable one-way puzzles. We show the following results about this variant
@@
    There exists a quantum oracle relative to which one-way puzzles exist but efficiently verifiable one-way puzzles do not.
@@
Note that our separation in fact separates efficiently verifiable one-way puzzles from pseudorandom state generators with linear output. But it is not hard to show that pseudorandom state generators with logarithmic output can be used to build efficiently verifiable one-way puzzles. This, our final separation also provides a barrier to length reeduction for pseudorandom state generators, answering an open question of [ALY23b].
@@ 2.1 A cleaner construction of EFI pairs from any one-way puzzle
For a $\epsilon$ one-way function, the KL-hardness parameter is $\delta=-\log\epsilon$. Thus, for a standard one-way function, $\delta=\omega(\log\lambda)$. But the techniques of [HRV10] apply whenever $\delta=\frac{1}{\mathsf{poly}(\lambda)}$, and so the techniques of [VZ12] work just as well for weak one-way functions. Thus, the same construction gives a PRG from any weak one-way functions.
@@ 2.2 Combiners and universal constructions
As shown originally by Levin [Lev87] and formalized in [${}^{+}$05], this useful fact is essentially a corollary of the fact that there exists robust combiners for one-way functions. That is, given any two one-way function candidates $f$ and $g$, there is a construction $h^{f,g}$ such that $h$ is one-way as long as one of $f$ or $g$ is one-way.
@@ 2.4 Relationships with other QCCC primitives
Note that minimality of a primitive is not very hard to achieve. As an example, any primitive which can be built unconditionally is minimal in the same sense as a one-way function. But, importantly, one-way functions are also useful. One-way functions can be used to construct a large class of important cryptographic primitives, often referred to as symmetric key primitives or Minicrypt. In particular, one-way functions can be used to build (classical) digital signatures, symmetric encryption protocols, and commitment schemes. Thus, if we want to treat one-way puzzles as a central primitive for QCCC cryptography, it seems important that one-way puzzles imply at least one more directly useful QCCC primitive.
Unfortunately, existing results in this direction are noticeably weaker. In particular, as far as the authors are aware, the only QCCC primitive for which a construction is known from one-way puzzles is non-uniform EFID pairs [KT24]. On the other hand, there are some quantum output implications of one-way puzzles. In particular, it is known how to build commitments with quantum output (and all equivalent primitives) from one-way puzzles [KT24].
@@ 2.5 Efficiently verifiable one-way puzzles
Theorem 8 (Theorems A.4 and A.6 from [KT24] and Theorems 15 and 16 in this paper).
@@
2.6 Random Input $\mathsf{OWPuzz}$
@@ 4.1 Definitions of QCCC primitives
    Opening stage: both parties receive the transcript $z$ as well as their private quantum states $\rho_{Com}$ and $\rho_{Rec}$ respectively. They then interact using only classical messages. At the end of the stage, the receiver either outputs a message or the reject symbol $\bot$.
@@ 5 Constructions of $\mathsf{EV-OWPuzz}$ from QCCC primitives
[…]
@@ 6 Efficiently verifiable one way puzzles can be broken with a $\mathsf{QCMA}$ oracle
For every efficiently verifiable one way puzzle $(\mathsf{Samp},\mathsf{Ver})$, there exists a promise problem $\Pi\in\mathsf{PromiseQCMA}$ and a QPT algorithm $\mathcal{A}^{\Pi}$ with oracle access to $\Pi$ which breaks security. That is
$\Pr_{\mathsf{Samp}(1^{\lambda})\to(k,s)}[\mathsf{Ver}(\mathcal{A}(s),s)\to 1]\geq\frac{1}{\mathsf{poly}(\lambda)}$
@@
    (yes): $(x,s)$ is a yes instance if there exists $y$ such that $\Pr[\mathsf{Ver}(x||y,s)\to 1]\geq\frac{2}{3}$
@@
    (no): $(x,s)$ is a no instance if for all $y$, $\Pr[\mathsf{Ver}(x||y,s)\to 1]<\frac{2}{3}$
@@
We will then show how to use oracle access to $\Pi$ to break the $\mathsf{EV-OWPuzz}$. This is just a standard search-to-decision reduction. The algorithm works as follows:
  1. 1.
    On input $s$, we define $k=k_{1},\dots,k_{n}$ bit by bit.
  2. 2.
    For $i=1,\dots,n$
    1. (a)
      Check if $((k_{1},\dots,k_{i-1},0),s)$ is a yes instance in $\Pi$. If so, set $k_{i}=0$.
    2. (b)
      Otherwise, check if $((k_{1},\dots,k_{i-1},1),s)$ is a yes instance in $\Pi$. If so, set $k_{i}=1$.
    3. (c)
      Otherwise, output $\bot$.
  3. 3.
    Output $s$.
Observe that as long as this algorithm does not output $\bot$, $(k,s)$ is a yes instance in $\Pi$ and so $\Pr[\mathsf{Ver}(k,s)\to 1]\geq\frac{2}{3}$.
[…]
Then, a simple induction argument says that as long as $s$ is a yes instance of $\Pi$, then $\mathcal{A}(s)$ always outputs a key $k$ such that $\Pr[\mathsf{Ver}(k,s)\to 1]\geq\frac{2}{3}$.
Correctness says that
$\Pr_{\mathsf{Samp}(1^{\lambda})\to(k,s)}[\mathsf{Ver}(k,s)\to 1]\geq 1-\mathsf{negl}(\lambda)$
@@
$\Pr_{\mathsf{Samp}(1^{\lambda})\to(k,s)}\left[\Pr[\mathsf{Ver}(k,s)\to 1]\geq\frac{2}{3}\right]\geq 1-\mathsf{negl}(\lambda)$
which means that with overwhelming probability, a sampled puzzle $s$ will be a yes instance of $\Pi$.
Putting this together, we get that
$\begin{split}\Pr_{\mathsf{Samp}(1^{\lambda})\to(k,s)}[\mathsf{Ver}(\mathcal{A}(s),s)\to 1]\\ \geq\Pr_{\mathsf{Samp}(1^{\lambda})\to(k,s)}[\mathsf{Ver}(\mathcal{A}(s),s)\to 1|\Pi(s)=1]\Pr_{\mathsf{Samp}(1^{\lambda})\to(k,s)}[\Pi(s)=1]\\ \geq\frac{2}{3}(1-\mathsf{negl}(\lambda))\\ \geq\frac{1}{\mathsf{poly}(\lambda)}\end{split}$
@@ 8.1 Amplification
Note that this construction only holds for normal one-way puzzles, as $\mathsf{Ver}^{\prime}$ is not efficient.
@@ 8.3 Universal Construction
Using standard padding arguments see [Lev87, HKNY23] we can show that if there exists a $(\alpha,\beta)$ $\mathsf{EV-OWPuzz}$ with both $\mathsf{Ver}$ and $\mathsf{Samp}$ running in time less than $\lambda^{c}$, then there exists a $(\alpha^{(1/c)},\beta^{(1/c)})$ $\mathsf{EV-OWPuzz}$ with $\mathsf{Ver}^{\prime}$ and $\mathsf{Samp}^{\prime}$ running in time less than $\lambda^{3}$. If $\mathsf{EV-OWPuzz}$’s exist then there exists a $(\mathsf{negl},\mathsf{negl})\mathsf{EV-OWPuzz}$ and therefore a $(\mathsf{negl},\mathsf{negl})\mathsf{EV-OWPuzz}$ running in time less than $\lambda^{3}$.
If some one-way puzzle exists then for some $i,j$, $(\mathsf{Samp}_{i},\mathsf{Ver}_{j})$ will be a one way puzzle with both algorithms running in time less than $n^{3}$. Once $\lambda\geq\max(i,j)$ a one way puzzle will one of the candidates. By Remark 4 we know that$(\mathsf{Samp}_{U},\mathsf{Ver}_{U})$ this will be a $\mathsf{OWPuzz}$. ∎
@@ 9.1 $\mathsf{OWPuzz}$s imply non-uniform EFID
An $\nu^{*}$-non-uniform EFID pair is a QPT algorithm $G_{\nu}(1^{\lambda},b)$ with classical parameter-dependent advice $\nu$. On input a unary security parameter $\lambda$ and $b\in\{0,1\}$ outputs a classical string such that
@@
This argument will follow from adapting the techniques of [VZ12] and [HRV10]. We recall a number of technical lemmas from these papers, and observe that since these lemmas rely only on black-box techniques, they also hold in the post-quantum setting. While we will focus on the non-uniform setting, the results should also hold against uniform adversaries by adapting the arguments of [VZ12] and [HRV10] in the uniform setting.

### v2

@@ Abstract
Recent work has introduced the “Quantum-Computation Classical-Communication” (QCCC) (Chung et. al.) setting for cryptography. There has been some evidence that One Way Puzzles ($\mathsf{OWPuzz}$) are the natural central cryptographic primitive for this setting (Khurana and Tomer). For a primitive to be considered central it should have several characteristics. It should be well behaved (which for this paper we will think of as having amplification, combiners, and universal constructions); it should be implied by a wide variety of other primitives; and it should be equivalent to some class of useful primitives. We present combiners, correctness and security amplification, and a universal construction for $\mathsf{OWPuzz}$. Our proof of security amplification uses a new and cleaner construction of EFI from $\mathsf{OWPuzz}$ (in comparison to the result of Khurana and Tomer) that generalizes to weak $\mathsf{OWPuzz}$ and is the most technically involved section of the paper. It was previously known that $\mathsf{OWPuzz}$ are implied by other primitives of interest including commitments, symmetric key encryption, one way state generators ($\mathsf{OWSG}$), and therefore pseudorandom states ($\mathsf{PRS}$). However we are able to rule out $\mathsf{OWPuzz}$’s equivalence to many of these primitives by showing a black box separation between general $\mathsf{OWPuzz}$ and a restricted class of $\mathsf{OWPuzz}$ (those with efficient verification, which we call $\mathsf{EV-OWPuzz}$). We then show that $\mathsf{EV-OWPuzz}$ are also implied by most of these primitives, which separates them from $\mathsf{OWPuzz}$ as well. This separation also separates extending $\mathsf{PRS}$ from highly compressing $\mathsf{PRS}$ answering an open question of Ananth et. al.
@@ 1 Introduction
In the realm of cryptography, there is perhaps no primitive more important than one-way functions. A one-way function is an efficiently computable deterministic function which is easy to compute, but hard to invert. Although at first glance the definition seems simple, one-way functions are special for several reasons. First and foremost, one-way functions are “minimal.” If modern cryptography exists in any form, then one-way functions must also exist [HILL99, IL89, Imp95]. Furthermore, pretty much all of these constructions are obvious. Second, one-way functions are “useful.” There is a large class of cryptographic primitives (known as Minicrypt) which can all be built from and are equivalent to one-way functions [Imp95]. Included in Minicrypt are symmetric key encryption, pseudorandom generators, and commitment schemes [HILL99, GGM86, Nao91]. Finally, one-way functions are “well-behaved.” They satisfy several natural properties [Lev87], and are equivalent to most of their variants [Yao82, IL89]. Due to these three characteristics of one-way functions, one of the most useful things to do when trying to understand a new classical cryptographic primitive is to compare it to a one-way function.
@@
In the classical setting, it appears that one-way functions serve as an effective minimal primitive. In the quantum output setting, EFI pairs are a promising candidate for our minimal primitive. A number of recent works have also considered a hybrid setting, primitives where the cryptographic algorithms are quantum, but all communication and outputs are classical [ACC+22, ALY23a]  [CLM23, KT24]. In the style of [ACC+22], we will refer to this as the quantum computation classical communication (QCCC) setting. An immediate and natural question about this setting is “what is a good central primitive?”
@@
  6. 6.
    We show that one-way puzzles are equivalent to ”distributional” one-way puzzles. For a distributional one-way puzzle, given a puzzle it is hard to sample from the distribution over keys.
We also consider in detail an important restricted variant of one-way puzzles, which was first introduced under the name “hard quantum planted problem for QCMA,” [KNY23], but which we will refer to as efficiently verifiable one-way puzzles. We show the following results about this variant
@@
    There exists a quantum oracle relative to which one-way puzzles and pseudorandom states exist but efficiently verifiable one-way puzzles do not.
@@
In addition to this, since pesudorandom states also exist under the same oracle, our results provide a barrier to building most QCCC primitives from pseudorandom states.
Note that our separation in fact separates efficiently verifiable one-way puzzles from pseudorandom state generators with linear output. But it is not hard to show that pseudorandom state generators with logarithmic output can be used to build efficiently verifiable one-way puzzles. Thus, our final separation also provides a barrier to length reduction for pseudorandom state generators, answering an open question of [ALY23b].
@@ 2.1 A cleaner construction of EFI pairs from any one-way puzzle
For an $\epsilon$ one-way function, the KL-hardness parameter is $\delta=-\log\epsilon$. Thus, for a standard one-way function, $\delta=\omega(\log\lambda)$. But the techniques of [HRV10] apply whenever $\delta=\frac{1}{\mathsf{poly}(\lambda)}$, and so the techniques of [VZ12] work just as well for weak one-way functions. Thus, the same construction gives a PRG from any weak one-way functions.
@@ 2.2 Combiners and universal constructions
As shown originally by Levin [Lev87] and formalized in [HKN+05], this useful fact is essentially a corollary of the fact that there exists robust combiners for one-way functions. That is, given any two one-way function candidates $f$ and $g$, there is a construction $h^{f,g}$ such that $h$ is one-way as long as one of $f$ or $g$ is one-way.
@@ 2.4 Relationships with other QCCC primitives
Note that minimality of a primitive is not very hard to achieve. As an example, any primitive which can be built unconditionally is at least as minimal as a one-way function. But, importantly, one-way functions are also useful. One-way functions can be used to construct a large class of important cryptographic primitives, often referred to as symmetric key or Minicrypt primitives. In particular, one-way functions can be used to build (classical) digital signatures, symmetric encryption protocols, and commitment schemes. Thus, if we want to treat one-way puzzles as a central primitive for QCCC cryptography, it seems important that one-way puzzles imply at least one more directly useful QCCC primitive.
Unfortunately, existing results in this direction are noticeably weaker. In particular, as far as the authors are aware, the only QCCC primitive for which a construction is known from one-way puzzles is non-uniform EFID pairs [KT24]. Instead, there are quantum output implications of one-way puzzles. In particular, it is known how to build commitments with quantum output (and all equivalent primitives) from one-way puzzles [KT24].
@@ 2.5 Efficiently verifiable one-way puzzles
Theorem 8 (Theorems A.4 and A.6 from [KT24] and Theorems 16 and 17 in this paper).
@@
2.6 Equivalence to variant definitions
Random input one way puzzles
@@ 2.6 Equivalence to variant definitions
Distributional one way puzzles
We can also consider an analogue of distributional one way functions, which are known to be equivalent to one way functions. That is, we say that a puzzle is distributionally one way if, given a puzzle, it is hard to sample the conditional distribution over keys. Note that every one way puzzle is a distributional one way puzzle. Interestingly, distributional one way puzzles do not need a verification algorithm, since their security comes from the sampled distribution.
We show that distributional one way puzzles are equivalent to one way puzzles. This follows directly from our techniques for amplification. In particular, a distributional one way puzzle is KL-hard to sample, and so can thus be used to build strong one way puzzles.
Theorem 14.
If there exists a distributional one-way puzzle, then there exists a one way puzzle.
We remark that as there is no verifier, there is not a natural definition of an ”efficiently verifiable” distributional one-way puzzle.
@@ 4.1 Definitions of QCCC primitives
    Correctness: For all sufficiently large $\lambda$, outputs of the sampler pass verification with overwhelming probability
@@
    Security: Given a puzzle $s$, it is computationally infeasible to find a key $s$ which verifies. That is, for all non-uniform QPT algorithms $\mathcal{A}$, for all sufficiently large $\lambda$,
@@
    (Non-interactive) opening stage: both parties receive the transcript $z$ as well as their private quantum states $\rho_{Com}$ and $\rho_{Rec}$ respectively. The committer $Com$ sends a single message $d$ to the receiver $Rec$. At the end of the stage, the receiver either outputs a message or the reject symbol $\bot$.
@@ 5 Constructions of $\mathsf{EV-OWPuzz}$ from QCCC primitives
    $\mathsf{Samp}$: Pick $m$ uniformly at random. Run $Com(m)\to(c,d)$. Output $(k=(m,d),s=c)$.
@@ 6 Efficiently verifiable one way puzzles can be broken with a $\mathsf{QCMA}$ oracle
Proposition 1 (From [INN+22, ABOBS22]).
There exists a search-to-decision reduction for $\mathsf{PromiseQCMA}$. Formally, there exists a promise problem $\Pi^{*}\in\mathsf{PromiseQCMA}$ such that for every $\Pi\in\mathsf{PromiseQCMA}$ with verifier $\mathcal{V}$, there exists a QPT algorithm $\mathcal{A}^{\Pi^{*}}$ such that for all $x$ such that $\Pi(x)=1$,
$\Pr[\Pr[\mathcal{V}(\mathcal{A}^{\Pi}(x),x)\to 1]\geq\frac{2}{3}]\geq\frac{1}{2}$
@@
For every efficiently verifiable one way puzzle $(\mathsf{Samp},\mathsf{Ver})$, there exists a promise problem $\Pi^{*}\in\mathsf{PromiseQCMA}$ and a QPT algorithm $\mathcal{A}^{\Pi^{*}}$ with oracle access to $\Pi^{*}$ which breaks security. That is
$\Pr_{\mathsf{Samp}(1^{\lambda})\to(k,s)}[\mathsf{Ver}(\mathcal{A}^{\Pi^{*}}(s),s)\to 1]\geq\frac{1}{\mathsf{poly}(\lambda)}$
@@
    (yes): $s$ is a yes instance if there exists $k$ such that $\Pr[\mathsf{Ver}(k,s)\to 1]\geq\frac{2}{3}$
@@
    (no): $s$ is a no instance if for all $k$, $\Pr[\mathsf{Ver}(k,s)\to 1]<\frac{1}{3}$
@@
Our promise problem $\Pi^{*}$ will be the problem from Proposition 1 and our algorithm $\mathcal{A}^{\Pi^{*}}$ will be the algorithm from Proposition 1 corresponding to $\Pi$. In particular, we have that for all $s$ such that there exists a $k$ with $\Pr[\mathsf{Ver}(k,s)\to 1]\geq\frac{2}{3}$, then
$\Pr[\Pr[\mathsf{Ver}(\mathcal{A}^{\Pi^{*}}(s),s)\to 1]\geq\frac{2}{3}]\geq\frac{1}{2}$
In particular, this means that
$\Pr[\mathsf{Ver}(\mathcal{A}^{\Pi^{*}}(s),s)\to 1]\geq\frac{1}{3}$
Correctness states that
$\Pr_{\mathsf{Samp}(1^{\lambda})\to(k,s)}[\mathsf{Ver}(k,s)\to 1]\geq 1-\mathsf{negl}(\lambda).$
@@
$\Pr_{\mathsf{Samp}(1^{\lambda})\to(k,s)}[\Pr[\mathsf{Ver}(k,s)\to 1]\geq\frac{2}{3}]\geq 1-\mathsf{negl}(\lambda)\geq\frac{1}{2}$
In particular, $s$ is a yes instance of $\Pi$ with all but negligible probability.
Putting things together, we get that
$\Pr_{\mathsf{Samp}(1^{\lambda})\to(k,s)}[\mathsf{Ver}(\mathcal{A}^{\Pi^{*}}(s),s)\to 1]\geq\frac{1}{6}$
and we are done. ∎
@@ 8.1 Amplification
Note that this construction only holds for standard inefficiently verifiable one-way puzzles, as $\mathsf{Ver}^{\prime}$ is not efficient.
@@ 8.3 Universal Construction
Using standard padding arguments (see [Lev87, HKNY23]) we can show that if there exists a $(\alpha,\beta)$ $\mathsf{EV-OWPuzz}$ with both $\mathsf{Ver}$ and $\mathsf{Samp}$ running in time less than $\lambda^{c}$, then there exists a $(\alpha^{(1/c)},\beta^{(1/c)})$ $\mathsf{EV-OWPuzz}$ with $\mathsf{Ver}^{\prime}$ and $\mathsf{Samp}^{\prime}$ running in time less than $\lambda^{3}$. If $\mathsf{EV-OWPuzz}$’s exist then there exists a $(\mathsf{negl},\mathsf{negl})\mathsf{EV-OWPuzz}$ and therefore a $(\mathsf{negl},\mathsf{negl})\mathsf{EV-OWPuzz}$ running in time less than $\lambda^{3}$.
If any one-way puzzle exists then for some $i,j$, $(\mathsf{Samp}_{i},\mathsf{Ver}_{j})$ will be a one way puzzle with both algorithms running in time less than $n^{3}$. Once $\lambda\geq\max(i,j)$ a one way puzzle will be one of the candidates. By Remark 4 we know that $(\mathsf{Samp}_{U},\mathsf{Ver}_{U})$ will be a $\mathsf{OWPuzz}$. ∎
@@ 9.1 $\mathsf{OWPuzz}$s imply non-uniform EFID
A $\nu^{*}$-non-uniform EFID pair is a QPT algorithm $G_{\nu}(1^{\lambda},b)$ with classical parameter-dependent advice $\nu$. On input a unary security parameter $\lambda$ and $b\in\{0,1\}$ outputs a classical string such that
@@
This argument will follow from adapting the techniques of [VZ12] and [HRV10]. We recall a number of technical lemmas from these papers, and observe that since these lemmas rely only on quantum friendly black-box techniques, they also hold in the QCCC setting. While we will focus on the non-uniform setting, the results should also hold against uniform adversaries by adapting the arguments of [VZ12] and [HRV10] in the uniform setting.